# An Adaptive Online Smoother with Closed-Form Solutions and Information-Theoretic Lag Selection for Conditional Gaussian Nonlinear Systems

### v1

@@ 3.1 Optimal Online Forward-In-Time Discrete Smoother
which is equivalent to (3.13), with the details of this equivalence being shown in Appendix B.
@@
         $\boldsymbol{\mu}_{\text{f}}^{n}\leftarrow\boldsymbol{\mu}_{\text{f}}^{n-1}+\left(\boldsymbol{\Lambda}^{\mathbf{y},n-1}\boldsymbol{\mu}_{\text{f}}^{n-1}+\mathbf{f}^{\mathbf{y},n-1}\right)\Delta t+\left(\mathbf{R}_{\text{f}}^{n-1}(\boldsymbol{\Lambda}^{\mathbf{x},n-1})^{\dagger}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1}\right)$
                  ×$\times\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1}\big{)}^{-1}\left(\mathbf{x}^{n}-\mathbf{x}^{n-1}-\left(\boldsymbol{\Lambda}^{\mathbf{x},n-1}\boldsymbol{\mu}_{\text{f}}^{n-1}+\mathbf{f}^{\mathbf{x},n-1}\right)\Delta t\right)$;
@@
                   $-\big{(}\mathbf{R}_{\text{f}}^{n-1}(\boldsymbol{\Lambda}^{\mathbf{x},n-1})^{\dagger}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1}\big{)}\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1}\big{)}^{-1}\big{(}\boldsymbol{\Lambda}^{\mathbf{x},n-1}\mathbf{R}_{\text{f}}^{n-1}+(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{n-1}\big{)}\Big{)}\Delta t$;
@@
         $\mathbf{b}^{n-1}\leftarrow\boldsymbol{\mu}_{\text{f}}^{n-1}-\mathbf{E}^{n-1}\left((\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},n-1}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{n-1}+\mathbf{f}^{\mathbf{y},n-1}\Delta t\right)$
                     $+\mathbf{F}^{n-1}\left(\mathbf{x}^{n}-\mathbf{x}^{n-1}-(\boldsymbol{\Lambda}^{\mathbf{x},n-1}\boldsymbol{\mu}_{\text{{f}}}^{n-1}+\mathbf{f}^{\mathbf{x},n-1})\Delta t\right)$;
@@ 3.2 Understanding How the Online Forward-In-Time Discrete Smoother Works
Figure 3.1 presents a schematic diagram that intuitively explains how the online smoother operates using forward filtering and backward smoothing estimates. Although the figure focuses on the online discrete smoother mean for brevity, the same logic (as well as subsequent analysi) applies to the smoother covariance. In the diagram, the first superscript index corresponds to rows, while the second corresponds to columns. The discussion centers on the last red dashed box (last column), representing the discrete smoother estimates at the current observation (the $n$-th observation) over all time instants $j=0,\ldots,n$. Each column (fixed second index) illustrates the backward smoothing process for a given batch realization $\{\mathbf{x}^{0},\ldots,\mathbf{x}^{n}\}$, involving a forward pass to compute the filter estimate (Theorem 2.2) and a backward pass to calculate the smoother estimate (Theorem 2.3), indicated by the red arrows. Each row (fixed first index) shows how the discrete smoother estimates at a fixed time $t_{j}$ are updated as new observations arrive sequentially, using the online forward-in-time smoother algorithm. Information flows onto $\boldsymbol{\mu}_{\text{s}}^{j,n}$ from the filter estimates $\boldsymbol{\mu}_{\text{f}}^{j}$ and $\boldsymbol{\mu}_{\text{f}}^{n}$. The former, $\boldsymbol{\mu}_{\text{f}}^{j}$, influences it through forward online-smoothing in its row, transitioning towards $\boldsymbol{\mu}_{\text{s}}^{j,n-1}$ and to the next columns using the online forward-in-time smoother (3.11). Through $\boldsymbol{\mu}_{\text{s}}^{j,n-1}$, the information in $\boldsymbol{\mu}_{\text{f}}^{n-1}$ is also captured via the classical backward smoother in its column. The latter, $\boldsymbol{\mu}_{\text{f}}^{n}$, affects it via the backward smoother (3.3) in its column, implicitly through $\boldsymbol{\mu}_{\text{s}}^{n-1,n}$ in (3.11). These observations via Figure 3.1 are explicitly and mathematically stated by the fact that the amount of updated information incurred at time step $t_{j}$, due to the new observation $\mathbf{x}^{n}$, is proportional to the update at $t_{n-1}$ via the matrix $\mathbf{D}^{j,n-2}$, since
$\boldsymbol{\mu}_{\text{s}}^{j,n}-\boldsymbol{\mu}_{\text{s}}^{j,n-1}=\mathbf{D}^{j,n-2}\left(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\right)\quad\text{and}\quad\boldsymbol{\mu}_{\text{f}}^{n-1}=\boldsymbol{\mu}_{\text{s}}^{n-1,n-1}.$
@@
It becomes apparent from (3.11)–(3.12) that to understand how the online smoother update works, as well as to understand how information dissipates through it, we need to study the mathematical properties of the update tensor given in (3.13), or compactly in (3.18), which in turn implies the thorough study of the auxiliary operator
@@
for $j=0,1,\ldots,n-2$. Based on the definitions of (2.6)–(2.7), it is easy to see that
@@
where $\prec$” is to be understood in the Loewner partial ordering sense over the convex cone of nonnegative-definite matrices, then by definition the spectral radius of $\mathbf{E}^{j}$ is less than $1$ uniformly over $j$,
@@
It can be shown that (3.19) is actually a necessary condition as to establish the mean-square stability of the smoother Gaussian statistics [77, 78, 79, 80, 81, 82], as can be clearly seen by the damping in the backward random differential equation that the smoother covariance satisfies in (2.9) of Theorem 2.3, with its discrete counterpart given in (3.4), and where the same damping coefficient also appears in the smoother mean’s evolution equation, since after some brief algebraic manipulations we have
$\overleftarrow{{\,\rm d}\boldsymbol{\mu}_{\text{{s}}}}=\left(-(\mathbf{A}+\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1})\boldsymbol{\mu}_{\text{{s}}}+\mathbf{f}^{\mathbf{y}}-\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1}\boldsymbol{\mu}_{\text{{f}}}\right){\,\rm d}t+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}\left(\overleftarrow{{\,\rm d}\mathbf{x}}+\mathbf{f}^{\mathbf{x}}{\,\rm d}t\right),$
with its discrete counterpart given in (3.3).
@@
Note that the filter covariance matrix $\mathbf{R}_{\text{f}}^{j}$ reflects uncertainties in both the observed and unobserved dynamics, as it is conditioned on observations. If the observed variables have much less uncertainty than the unobserved process, the eigenvalues of $(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{j}$ will be much larger than those of $\mathbf{R}_{\text{f}}^{j}$, as suggested by discretizing the random Riccati equation (2.4), since the quadratic term is weighted by $\boldsymbol{\Gamma}:=(\boldsymbol{\Lambda}^{\mathbf{x}})^{\dagger}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}\boldsymbol{\Lambda}^{\mathbf{x}}$ which is being taken out from the nonnegative-definite stochastic forcing $\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}}$. Consequently, by this reasoning and (3.5), the spectrum of $\mathbf{E}^{j}$ should concentrate near the origin and within the unit disk in the complex plane. In this scenario, observations become highly informative. They provide sufficient information to update the estimated state over a short time interval. In contrast, their influence on earlier times decays exponentially, which reflects the exponential mean-square stability of the smoother covariance equations. Numerical examples, such as the nonlinear dyad-interaction model in Section 4.1, show that even with intermittent instabilities, extreme events, and noise cross-interactions, the spectral radius of the update matrix $\mathbf{D}^{j,n-2}$ exhibits exponential decay backwards in time, with the rate increasing under the observance of extreme events in the observed time series; observing the intermittent extreme events increases the signal-to-noise ratio, thus reducing the observational uncertainty. This also leads to larger required lag values in the adaptive-lag strategy for the optimal online forward-in-time smoother discussed in Section 3.4.2, as to effectively resolve the intermittent instabilities at their onset.
@@
To make things more concrete, we show that in the case of a simple two-dimensional linear Gaussian model, the value of $\mathbf{E}^{j}\equiv E^{j}$ (which in this case is just a scalar), is necessarily in $(-1,1)$, thus concretely establishing (3.20) and by extension, due to the one-dimensional observable and unobserved state spaces, leads to $\mathbf{D}^{j,n-2}\equiv D^{j,n-2}\in(-1,1)$ for this specific example. Consider the following two-dimensional linear Gaussian model:
${\,\rm d}x$ $=\left(a_{11}x+a_{12}y\right)+\sigma_{1}{\,\rm d}W_{1},$
${\,\rm d}y$ $=\left(a_{21}x+a_{22}y\right)+\sigma_{2}{\,\rm d}W_{2},$
@@
where $a_{11}$, $a_{12}$, $a_{21}$, $a_{22}\in\mathbb{R}$, $\sigma_{1}>0$, and $\sigma_{2}>0$ are constants, and $x$ and $y$ are the observed and unobserved variables, respectively. Note that we assume the absence of noise cross-interaction. Here, due to the linearity of the system, $a_{22}$ needs to be negative in order to guarantee the existence of the statistical equilibrium state of the coupled system when $y$ is assumed to be the unobservable (this also becomes apparent from (3.21) later on) [75, 83, 84, 78]. Since the coefficients are constants, the equilibrium solution for the filter covariance, $R_{\text{f}}^{\text{eq}}$, can thus be solved via the following steady state Riccati equation (see (2.4)) [42, 47, 75]:
$2a_{22}R_{\text{f}}^{\text{eq}}+\sigma_{2}^{2}=\frac{(R_{\text{f}}^{\text{eq}}a_{12})^{2}}{\sigma_{1}^{2}},$ (3.21)
@@
$R_{\text{f}}^{\text{eq}}=\frac{a_{22}\sigma_{1}^{2}+\sigma_{1}\sqrt{\Psi}}{a_{12}^{2}},$ (3.22)
where $\Psi=a_{22}^{2}\sigma_{1}^{2}+a_{12}^{2}\sigma_{2}^{2}$. Observe now that by plugging-in (3.22) into the expression found in (3.19),
@@
$G^{y}$ $=a_{22}+\frac{\sigma_{2}^{2}}{R_{\text{f}}^{\text{eq}}}=\frac{1}{R_{\text{f}}^{\text{eq}}}\left(a_{22}\frac{a_{22}\sigma_{1}^{2}+\sigma_{1}\sqrt{\Psi}}{a_{12}^{2}}+\sigma_{2}^{2}\right)$
$=\frac{1}{R_{\text{f}}^{\text{eq}}}\left(\frac{a_{22}^{2}\sigma_{1}^{2}+a_{22}\sigma_{1}\sqrt{\Psi}+a_{12}^{2}\sigma_{2}^{2}}{a_{12}^{2}}\right)=\frac{\sqrt{\Psi}}{a_{12}^{2}R_{\text{f}}^{\text{eq}}}\left(\sqrt{\Psi}+a_{22}\sigma_{1}\right)>0,$
where we have used the definition of $\Psi$, the fact that $\sqrt{\Psi}+a_{22}\sigma_{1}=\sqrt{a_{22}^{2}\sigma_{1}^{2}+a_{12}^{2}\sigma_{2}^{2}}+a_{22}\sigma_{1}>0$, and that $R_{\text{f}}^{\text{eq}}>0$. This validates condition (3.19), and so the auxiliary constant $E^{j}$ is always within the interval $(-1,1)$, meaning (3.20) is satisfied. By extension, since everything is just a scalar, this implies $\mathbf{D}^{j,n-2}\equiv D^{j,n-2}\in(-1,1)$ because of (3.18) (see Appendix B).
In general, except in the trivial cases where the unobserved state space is one-dimensional, the condition (3.20) on the constituents of the product defining the update matrix $\mathbf{D}^{j,n-2}$ in (3.18) do not guarantee that the update tensor $\mathbf{D}^{j,n-2}$ itself will also have a spectral radius less than $1$. This is because, generally, the spectral radius of a product is not dominated by the product of the associated spectral radii unless rather restrictive conditions are applied, with sufficient conditions including simultaneous triangularization of the factors over $\mathbb{C}$ [85] (e.g., the factors commute, as a consequence of Gelfand’s theorem) or the factors being radial matrices [86]. Consequently, predicting how each observation will affect state estimation through online smoother updates can be complex and highly dependent on the specific structure of the CGNS. This variability underscores the need for a simple numerical procedure to assess the impact region of each new observation, as described by the recursive backward difference equations for the discrete smoother posterior Gaussian statistics in (3.11)–(3.12), which will in turn aid in constructing an effective adaptive lag strategy for the online smoother (see Section 3.4). Nevertheless, even in the ideal setting where control on the spectrum of $\mathbf{D}^{j,n-2}$ can be established, an online smoother which defines its adaptive lag based on the eigenvalues of the update matrix will incur computational costs in the order of $O(l^{3})$, which for high-dimensional systems is inexcusable.
As the system dimension and observational period length increase, the computational overhead for updating the online smoothing procedure becomes significant, particularly due to the storage of auxiliary matrices like $\mathbf{E}^{j}$, which are essential for calculating the update matrix $\mathbf{D}^{j,n-2}$, as well as the filter and smoother covariance matrices. Additionally, constant recalculation of the covariance tensor can lead to $O(l^{3})$ computations per update, especially when $k\approx l$. Given that the impact of new observations on past states decays exponentially (linked to the aforementioned spectral properties of $\mathbf{D}^{j,n-2}$ and $\mathbf{E}^{j}$), it is advantageous to develop a fixed- or adaptive-lag online smoother [17, 27, 43, 44]. In such a way, the impact region of each observation is either finite and predetermined or computable based on the dynamical and statistical properties of the underlying system.

### v2

@@ 3.1 Optimal Online Forward-In-Time Discrete Smoother
which is equivalent to (3.12), with the details of this equivalence being shown in Appendix B. (The curved arrow pointing to the right above the product symbol in (3.17) indicates the order or direction with which we expand said product.)
@@
    $\boldsymbol{\mu}_{\text{f}}^{n}\leftarrow\boldsymbol{\mu}_{\text{f}}^{n-1}+(\boldsymbol{\Lambda}^{\mathbf{y},n-1}\boldsymbol{\mu}_{\text{f}}^{n-1}+\mathbf{f}^{\mathbf{y},n-1})\Delta t+(\mathbf{R}_{\text{f}}^{n-1}(\boldsymbol{\Lambda}^{\mathbf{x},n-1})^{\dagger}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1})$
       $\times\big((\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1}\big)^{-1}\big(\mathbf{x}^{n}-\mathbf{x}^{n-1}-(\boldsymbol{\Lambda}^{\mathbf{x},n-1}\boldsymbol{\mu}_{\text{f}}^{n-1}+\mathbf{f}^{\mathbf{x},n-1})\Delta t\big)$;
@@
        $(\mathbf{R}_{\text{f}}^{n-1}(\boldsymbol{\Lambda}^{\mathbf{x},n-1})^{\dagger}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1})\big((\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{n-1}\big)^{-1}(\boldsymbol{\Lambda}^{\mathbf{x},n-1}\mathbf{R}_{\text{f}}^{n-1}+(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{n-1})\big)\Delta t$;
@@
    $\mathbf{b}^{n-1}\leftarrow\boldsymbol{\mu}_{\text{f}}^{n-1}-\mathbf{E}^{n-1}\big((\mathbf{I}_{l\times l}+\boldsymbol{\Lambda}^{\mathbf{y},n-1}\Delta t)\boldsymbol{\mu}_{\text{{f}}}^{n-1}+\mathbf{f}^{\mathbf{y},n-1}\Delta t\big)$
        $\mathbf{F}^{n-1}\big(\mathbf{x}^{n}-\mathbf{x}^{n-1}-(\boldsymbol{\Lambda}^{\mathbf{x},n-1}\boldsymbol{\mu}_{\text{{f}}}^{n-1}+\mathbf{f}^{\mathbf{x},n-1})\Delta t\big)$;
@@ 3.2 Understanding How the Online Forward-In-Time Discrete Smoother Works
Figure 3.1 presents a schematic diagram that intuitively explains how the online smoother operates using forward filtering and backward smoothing estimates. Although the figure focuses on the online discrete smoother mean for brevity, the same logic (as well as subsequent analysis) applies to the smoother covariance. In the diagram, the first superscript index corresponds to rows, while the second corresponds to columns. The discussion centers on the last red dashed box (last column), representing the discrete smoother estimates at the current observation (the $n$-th observation) over all time instants $j=0,\ldots,n$. Each column (fixed second index) illustrates the backward smoothing process for a given batch realization $\{\mathbf{x}^{0},\ldots,\mathbf{x}^{n}\}$, involving a forward pass to compute the filter estimate (Theorem 2.2) and a backward pass to calculate the smoother estimate (Theorem 2.3), indicated by the red arrows. Each row (fixed first index) shows how the discrete online smoother estimates at a fixed time $t_{j}$ are updated as new observations arrive sequentially, using the online forward-in-time smoother algorithm. Information flows into $\boldsymbol{\mu}_{\text{s}}^{j,n}$ from the filter estimates $\boldsymbol{\mu}_{\text{f}}^{j}$ and $\boldsymbol{\mu}_{\text{f}}^{n}$. The former, $\boldsymbol{\mu}_{\text{f}}^{j}$, influences it through forward online-smoothing in its row, transitioning towards $\boldsymbol{\mu}_{\text{s}}^{j,n-1}$ and to the next columns using the online forward-in-time smoother (3.10). (Through $\boldsymbol{\mu}_{\text{s}}^{j,n-1}$, the information in $\boldsymbol{\mu}_{\text{f}}^{n-1}$ is also captured via the classical backward smoother in its column.) The latter, $\boldsymbol{\mu}_{\text{f}}^{n}$, affects it via the backward smoother (3.3) in its column, implicitly through $\boldsymbol{\mu}_{\text{s}}^{n-1,n}$ in (3.10). These deductions, stemming from Figure 3.1, are explicitly and mathematically expressed by the fact that the amount of updated information incurred at time step $t_{j}$, due to the new observation $\mathbf{x}^{n}$, is proportional to the update at $t_{n-1}$ via the matrix $\mathbf{D}^{j,n-2}$, since
$\boldsymbol{\mu}_{\text{s}}^{j,n}-\boldsymbol{\mu}_{\text{s}}^{j,n-1}=\mathbf{D}^{j,n-2}\big(\boldsymbol{\mu}_{\text{s}}^{n-1,n}-\boldsymbol{\mu}_{\text{f}}^{n-1}\big)\quad\text{and}\quad\boldsymbol{\mu}_{\text{f}}^{n-1}=\boldsymbol{\mu}_{\text{s}}^{n-1,n-1}.$
@@
3.2.1 How the Online Forward-In-Time Discrete Smoother’s Mechanisms Guide the Development of an Adaptive-Lag Algorithm
It becomes apparent from (3.10)–(3.11) that to understand how the online smoother update works, as well as to figure out how information dissipates through it, we need to study the mathematical properties of the update tensor given in (3.12), or compactly in (3.17), which in turn implies the thorough study of the auxiliary operator
@@
for $j=0,1,\ldots,n-2$. Such an analysis will also assist in the formulation of an adaptive lag variant for the online discrete smoother, that implicitly optimizes the use of data for the reduction of the computational overhead.
Based on the definitions of (2.6)–(2.7), it is easy to see that
@@
where $\prec$” is to be understood in the Loewner partial ordering sense over the convex cone of nonnegative-definite matrices, then by definition the spectral radius of $\mathbf{E}^{j}$ is less than $1$ uniformly over $j$:
@@
Notably, (3.18) is a necessary condition to establish mean-square stability of the smoother Gaussian statistics [63, 64, 65, 66], as can be clearly seen by the damping in the backward random differential equation that the smoother covariance satisfies in (2.9) of Theorem 2.3, with its discrete counterpart given in (3.4), and where the same damping coefficient also appears in the smoother mean’s linear evolution equation, since after a few algebraic manipulations we have
$\overleftarrow{{\,\rm d}\boldsymbol{\mu}_{\text{{s}}}}=\big(-(\mathbf{A}+\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1})\boldsymbol{\mu}_{\text{{s}}}+\mathbf{f}^{\mathbf{y}}-\mathbf{B}\mathbf{R}_{\text{{f}}}^{-1}\boldsymbol{\mu}_{\text{{f}}}\big){\,\rm d}t+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}\big(\overleftarrow{{\,\rm d}\mathbf{x}}+\mathbf{f}^{\mathbf{x}}{\,\rm d}t\big),$
@@
with its discrete counterpart given in (3.3). Probing into (3.19), note that the filter covariance matrix, $\mathbf{R}_{\text{f}}^{j}$, reflects uncertainties in both the unobserved and observed dynamics, as it is conditioned on the current observable $\sigma$-algebra. If the observed variables have much less uncertainty than the unobserved process, the eigenvalues of $(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})^{j}$ will be much larger than those of $\mathbf{R}_{\text{f}}^{j}$, as suggested by discretizing the random Riccati equation in (2.4), since the quadratic term is weighted by $\boldsymbol{\Gamma}:=(\boldsymbol{\Lambda}^{\mathbf{x}})^{\dagger}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}\boldsymbol{\Lambda}^{\mathbf{x}}$ and then subtracted from the nonnegative-definite stochastic forcing $\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}}$ (in the update of the forecast uncertainty). Consequently, by this reasoning and (3.5), the spectrum of $\mathbf{E}^{j}$ should concentrate near the origin and within the unit disk in the complex plane, thus satisfying (3.19). Recalling that the control on the spectrum of $\mathbf{E}^{j}$ is determined by the exponential mean-square stability of the smoother posterior statistics, in this scenario observations become highly informative; they provide sufficient information to update the estimated state over a short time interval, with an exponentially-fast decaying impact at distant past time instants.
This is further reflected in the spectral radius of the update matrix, $\mathbf{D}^{j,n-2}$, which explicitly determines the influence of the observations on the discrete smoother updates during the recursive backward difference equations in (3.10)–(3.11). Numerical examples, such as the nonlinear dyad-interaction model in Section 4.1, show that even with intermittent instabilities and noise cross-interactions it exhibits an overall exponential decay backwards in time (i.e., in $j$ for each $n$), with an accelerating rate under the observance of extreme events in the observed time series. Therefore, any potent adaptive-lag strategy accompanying the forward-in-time online smoother in (3.10)–(3.11) should:
  • (a)
    Yield larger lag values during the generation of extreme events, where observations significantly inform and contribute to the smoother update, thus quickly resolving the intermittent instabilities. During this observational period (i.e., for these $n$), the spectral radius of $\mathbf{D}^{j,n-2}$ remains significant (close to $1$) for time instants (i.e., for $j$) which cover the initiation period of the extreme event.
  • (b)
    Produce smaller lags at periods of large signal-to-noise ratios in the observed variables, where the prior state estimates from the preceding observations are sufficiently skillful (e.g., during the demise of extreme events). During this observational period (i.e., for these $n$), the spectral radius of $\mathbf{D}^{j,n-2}$ showcases remarkable rapid exponential decay at time instants (i.e., for $j$) prior to and during the extreme event.
See Panels (g)–(i) of Figure 4.2 and Panels (a)–(c) of Figure 4.5 for numerical examples.
In general, except in trivial cases where the unobserved state space is one-dimensional (e.g., see Section 3.2.2), condition (3.19) on the constituents of the product defining the update matrix $\mathbf{D}^{j,n-2}$ in (3.17) does not guarantee that the update tensor $\mathbf{D}^{j,n-2}$ itself will also have a subunit spectral radius for $j$ sufficiently far back from n1n-1. (Such a condition is necessary for establishing rigorous mathematical results of convergence for (3.10)–(3.11), and reflects the decaying impact region of $\mathbf{x}^{n}$ on $\mathbf{y}^{j}$’s online smoother state estimation.) This is because, generally, the spectral radius of a product is not dominated by the product of the associated spectral radii unless rather restrictive conditions are assumed, with sufficient conditions including simultaneous triangularization of the factors over $\mathbb{C}$ [67] (e.g., the factors commute, as a consequence of Gelfand’s theorem) or the factors being radial matrices [68]. Consequently, predicting how each observation affects state estimation through the online smoother updates can be complex and highly dependent on the specific structure of the CGNS. This variability underscores the need for a simple numerical procedure to assess the impact region of each new observation, as described by the recursive backward difference equations for the discrete smoother posterior Gaussian statistics in (3.10)–(3.11), which in turn aids in constructing an effective adaptive lag strategy for the online smoother. Nevertheless, even in the ideal setting where control on the spectrum of $\mathbf{D}^{j,n-2}$ can be established, an online smoother which defines its adaptive lag based on the eigenvalues of the update matrix incurs computational costs in the order of $O(l^{3})$, which for high-dimensional systems is inexcusable. This is why, in Section 3.4, we develop an information theory-based adaptive-lag approach to the online discrete smoother which efficiently approximates the observational impact region, while implicitly and effectively capturing the essence of the exact method defined by the spectral properties of the update matrix $\mathbf{D}^{j,n-2}$.
As the system dimension and $n$ increase, the computational overhead for updating the online smoother becomes significant, particularly due to storing the auxiliary matrices needed for calculating the update matrix $\mathbf{D}^{j,n-2}$, specifically the $\mathbf{E}^{j}$ matrices, as well as the filter and online smoother covariance matrices. Additionally, constant recalculation of the covariance tensor can lead to $O(l^{3})$ computations per update, especially when $k\approx l$. Given that the impact of new observations on past states decays exponentially (linked to the aforementioned spectral properties of $\mathbf{D}^{j,n-2}$ and $\mathbf{E}^{j}$), it is advantageous to develop an adaptive-lag (or fixed-lag) online smoother [24, 38, 39], where the impact region of or lag at each observation is adaptively computed (or finite and predetermined) based on the dynamical and statistical properties of the CGNS.
3.2.2 Behavior of the Online Forward-In-Time Discrete Smoother for a Two-Dimensional Linear System with Additive Noise
@@
To make things more concrete, we show that in the case of a simple two-dimensional linear Gaussian model, the value of $\mathbf{E}^{j}\equiv E^{j}$ (which in this case is just a scalar), is necessarily in (1,1)(-1,1), thus concretely establishing (3.19) and by extension, due to the one-dimensional observable and unobserved state spaces, leads to $\mathbf{D}^{j,n-2}\equiv D^{j,n-2}\in(-1,1)$ for this specific example. Consider the following two-dimensional linear Gaussian model with additive noise:
${\,\rm d}x$ $=\big(\lambda_{x}y+f_{x}(t,x)\big)+\sigma_{x}{\,\rm d}W_{1},$
${\,\rm d}y$ $=\big(\lambda_{y}y+f_{y}(t,x)\big)+\sigma_{y}{\,\rm d}W_{2},$
@@
where $\lambda_{x}\in\mathbb{R}$, $\lambda_{y}<0$, $\sigma_{x}>0$, and $\sigma_{y}>0$ are constants and $f_{x}(t,x)$ is such that the $x$-dynamics are stable. As before, $x$ and $y$ are the observed and unobserved variables, respectively. Note that we assume the absence of noise cross-interaction, without loss of generality. Here, due to the linearity of the system, $\lambda_{y}$ needs to be negative in order to guarantee the existence of the statistical equilibrium state of the coupled system when $y$ is assumed to be the unobservable (this also becomes apparent from (3.20) later on) [62, 63, 69]. Since $\lambda_{x}$, $\lambda_{y}$ and the noise feedbacks are constant, the equilibrium solution for the filter covariance, $R_{\text{f}}^{\text{eq}}$, can thus be solved via the following steady state Riccati equation (see (2.4)) [41, 42, 62]:
$2\lambda_{y}R_{\text{f}}^{\text{eq}}+\sigma_{y}^{2}=\frac{(R_{\text{f}}^{\text{eq}}\lambda_{x})^{2}}{\sigma_{x}^{2}},$ (3.20)
@@
$R_{\text{f}}^{\text{eq}}=\frac{\lambda_{y}\sigma_{x}^{2}+\sigma_{x}\sqrt{\Psi}}{\lambda_{x}^{2}},$ (3.21)
where $\Psi=\lambda_{y}^{2}\sigma_{x}^{2}+\lambda_{x}^{2}\sigma_{y}^{2}$. Observe now that by plugging-in (3.21) into the expression found in (3.18),
@@
$G^{y}$ $=\lambda_{y}+\frac{\sigma_{y}^{2}}{R_{\text{f}}^{\text{eq}}}=\frac{1}{R_{\text{f}}^{\text{eq}}}\left(\lambda_{y}\frac{\lambda_{y}\sigma_{x}^{2}+\sigma_{x}\sqrt{\Psi}}{\lambda_{x}^{2}}+\sigma_{y}^{2}\right)$
$=\frac{1}{R_{\text{f}}^{\text{eq}}}\left(\frac{\lambda_{y}^{2}\sigma_{x}^{2}+\lambda_{y}\sigma_{x}\sqrt{\Psi}+\lambda_{x}^{2}\sigma_{y}^{2}}{\lambda_{x}^{2}}\right)=\frac{\sqrt{\Psi}}{\lambda_{x}^{2}R_{\text{f}}^{\text{eq}}}\big(\sqrt{\Psi}+\lambda_{y}\sigma_{x}\big)>0,$
where we have used the definition of $\Psi$, the fact that $\sqrt{\Psi}+\lambda_{y}\sigma_{x}=\sqrt{\lambda_{y}^{2}\sigma_{x}^{2}+\lambda_{x}^{2}\sigma_{y}^{2}}+\lambda_{y}\sigma_{x}>0$, and that $R_{\text{f}}^{\text{eq}}>0$. This validates condition (3.18), for sufficiently small step size $\Delta t$, and so the auxiliary constant $E^{j}$ is always within the interval (1,1)(-1,1), meaning (3.19) is satisfied. By extension, since everything is just a scalar, this implies $\mathbf{D}^{j,n-2}\equiv D^{j,n-2}\in(-1,1)$ because of (3.17) (see Appendix B).